# Improved Prediction Of Ligand-Protein Binding Affinities By Meta-Modeling Ho-Joon Lee1,2,*#, Prashant S. Emani3,*# and Mark B. Gerstein3,4,5,6,7#

### v3

@@ Discussion
The framework we present here, inspired by stacked generalization and super learner 74, 75, is based on the central concept that different models designed and trained for the same prediction task often carry model-specific biases. Hence, simple approaches to combining the results of their independent predictions may yield superior results if the biases are even partially independent and the models are complementary to some degree. Indeed, we find this to be borne out in the ligand-protein binding affinity prediction in several ways. First, we find that combining the predictions from empirical docking methods with those from the pre-trained or de novo-trained deep learning tools can yield gains in predictive power over the individual approaches (ED1-3 vs. E or D1-2). Second, utilizing models trained on the BindingDB as the base and refining them on the PDBbind resulted in significant improvements (D1 and D2 vs. D1F and D2F, respectively), demonstrating that fine-tuning pre-trained models could retain good predictive quality on both primary (BindingDB) and secondary (PDBbind) datasets (Supplementary Fig. S4). Third, even within the suite of deep learning tools, transforming their prediction scores through principal component analysis could improve the performance on the test set (i.e., principal component regression). This may be due to a de-noising of individual DL scores. Finally, in addition to performance improvement or synergy by the meta-models over the docking and DL models for a subset of complexes, the use of meta-models by canonical machine learning enables us to explicitly add features that may be associated with prediction deviations in the individual models. For example, we found that, for both empirical docking tools and some deep learning models, the deviations of the predictions from the experimental affinities were correlated with molecular weight. Explicitly adding in molecular weight as a meta-model feature resulted in the deviations of the final predictions being much less correlated with molecular weight.
@@
Overall, our framework corroborates existing work on ensemble learning, stacked generalization, super learner, and consensus-based improvements in predictions. Our machine-learning-based meta-models allow for the flexibility of adding or removing any component tools, physicochemical properties, or molecular descriptors as part of feature engineering. Models trained on multiple datasets (e.g. BindingDB and PDBbind) and on multiple data types (e.g. sequence and/or structure) may be unified to reduce individual biases and improve predictions. Furthermore, we posit that the DL models from DeepPurpose considered herein, being sequence-based, offer better scalability. Given that only ligand and protein sequences are required for the prediction, and not all-atom 3D structures, we expect that increasingly large databases could be used for training in future iterations by running high-throughput experimental binding affinity assays (without the need for concurrent X-ray or NMR experiments) as a filtering strategy in early drug discovery. Considering that the observed performances of these models for both benchmarks of CASF-2016 and the small-scale virtual screening are better or on par with more sophisticated structure-based DL models, there is an argument to be made in favor of more scalable approaches. Additionally, we note that the public and private experimental binding affinity databases are highly fragmented, in that they are produced separately, with potentially different types of assays, or at different scales. While efforts such as the PDBbind database manage to carefully curate a high-quality single dataset, not all data sources can be incorporated because of the type of data (e.g. experimental data without associated structures being available) or because of data access (e.g. behind a private paywall). Our meta-modeling may address some of these issues by pooling only prediction outputs from models trained on diverse public or private databases with or without available structures, protecting the databases themselves if necessary and maintaining performance quality. Subsequent in-depth analysis could also identify prediction synergy and complementarity of all available tools as we demonstrated in Fig. 6. Finally, there might be value in expanding families of meta-models with different input data, each family being optimized to target a particular protein or class of proteins and/or ligands as demonstrated in Tables 4 and 5.
@@
We acknowledge some limitations of our study. First, we primarily focused on a single dataset, PDBbind RefinedSet, for both training and testing. In fact, in light of recent positive results from training models on the full PDBbind General Set of ~19,000 complexes 42, 47, there is a scope for further scale-up of the model training process with larger data. It is worth mentioning, however, that when Jones et al. used the PDBbind general or refined set for model training, they observed the tradeoffs between data quality and quantity in performance of different models: SG-CNN performed better with the general set (in favor of data quantity); 3D-CNN performed better with the refined set (in favor of data quality). On the other hand, we started with DL models pre-trained on BindingDB of ~66,000 complexes and either employed them “out of the box” or fine-tuned them on the PDBbind RefinedSet (excluding the CoreSet). The two BindingDB-based meta-models, ED1 and ED2, perform better than any of the stand-alone docking or DL tools (Pearson correlation 0.62-0.64 vs. 0.47-0.59). We also tested the two PDBbind fine-tuned models, D1F and D2F, on the primary training set of BindingDB, to ascertain the degree to which the models retained their “out of the box” performance on the primary training set. We found reasonably good correlations for D1F and D2F (r = 0.636 to 0.711) (Supplementary Fig. S4) compared to the pre-trained D1 (r = 0.8) 57. However, the PDBbind-trained D3 showed poor prediction on BindingDB (r < 0.2). More importantly, while the CASF-2016 benchmark set has been well accepted, more benchmark sets of high quality would be needed to test and compare models in a more objective manner and also to challenge model generalization. Our virtual screening benchmarks are also very limited and our models tend to perform well for ligands with low molecular weight. In particular, the binary classification problem for screening of active or functional ligands would require a new model development for higher accuracy rather than repurposing exclusively affinity-based regression models as in this study. In this sense, the good performances of our fine-tuned DL models or meta-models do not imply model generalizability. We also note that we neither fine-tuned the docking tools nor conducted hyperparameter optimization for the DL models, which might help performance as well.
@@
Our proof-of-concept study of meta-modeling can be readily extended to integrate with (i.e., to add more meta features of) predictions from any other tools, such as HAC-Net, FAST, KDeep, and/or PerSpect ML, molecular properties or embeddings, or even predictions from the same base models trained on different orthogonal datasets, to address model bias and robustness, which is the key conceptual advantage of any ensemble or fusion model. In parallel to this type of horizontal extension (i.e., feature engineering), it is conceivable that meta-models may be expanded or stacked vertically, leading to deep meta-models. Our meta-modeling approach offers promising and flexible strategies for future model iterations to further improve ligand-protein binding affinity prediction.
@@
The BindingDB and PDBbind databases are publicly available. The code is available on our GitHub page at https://github.com/Lee1701/Lee2023a.
@@ Acknowledgements
1Dept. of Genetics and 2Yale Center for Genome Analysis, Yale University, New Haven, CT 06510, USA 3Dept. of Molecular Biophysics & Biochemistry, Yale University, New Haven, CT 06520, USA 4Program in Computational Biology & Bioinformatics, 5Dept. of Computer Science, 6Dept. of Statistics & Data Science, and 7Dept. of Biomedical Informatics & Data Science, Yale University, New Haven, CT 06520, USA
@@
Preparation of the receptor structure: The receptor structure is read in as a PDB file and preprocessed in PyMol (https://pymol.org). We used the pymol class and associated functions within our pipeline script. The preprocessing consists of the removal of all solvent and heteroatoms in the molecules. The resulting PDB file is then converted into the requisite PDBQT format using OpenBabel (http://openbabel.org). We used the obabel flags ‘-e = continue after errors’, ‘-p = add hydrogens appropriate for the pH’, and ‘–partialcharge gasteiger’ to calculate Gasteiger partial charges. The pH was set at 7.0. We included the PDBQT-specific flags ‘r = output as a rigid molecular’ and ‘p = Preserve atom indices from input file’. The implementation of this is written as:
@@
For the LIT-PCBA virtual screening exercise, we utilized balanced sets of 10 active and 10 inactive ligands randomly drawn from the combined training and validation cohorts of the LIT_PCBA database8. To run the docking tools on the datasets, we generated three-dimensional SDF structures for each of the ligands from their SMILES strings using the following command:
@@
[margin-bottom=0.0063in,margin-top=0mm,margin-right=0mm,margin-left=0mm,border-style=none,padding=0mm]![[Uncaptioned image]](extracted/5604664/a0000-img017.png)

### v4

@@ Discussion
The framework we present here, inspired by stacked generalization and super learner 74, 75, is based on the central concept that different models designed and trained for the same prediction task often carry model-specific biases. Hence, simple approaches to combining the results of their independent predictions may yield superior results if the biases are even partially independent and the models are complementary to some degree. Indeed, we find this to be borne out in the ligand-protein binding affinity prediction in several ways. First, we find that combining the predictions from empirical docking methods with those from the pre-trained or de novo-trained deep learning tools can yield gains in predictive power over the individual approaches (ED1-3 vs. E or D1-2). Second, utilizing models trained on the BindingDB as the base and refining them on the PDBbind resulted in significant improvements (D1 and D2 vs. D1F and D2F, respectively), demonstrating that fine-tuning pre-trained models could retain good predictive quality on both primary (BindingDB) and secondary (PDBbind) datasets (Supplementary Fig. S4). Third, even within the suite of deep learning tools, transforming their prediction scores through principal component analysis could improve the performance on the test set (i.e., principal component regression). This may be due to a de-noising of individual DL scores. Finally, in addition to performance improvement or synergy by the meta-models over the docking and DL models for a subset of complexes, the use of meta-models by canonical machine learning enables us to explicitly add features that may be associated with prediction deviations from the experimental affinities in the individual models. For example, we found that, for both empirical docking tools and some deep learning models, the deviations of the predictions from the experimental affinities were correlated with molecular weight. Explicitly adding in molecular weight as a meta-model feature resulted in the deviations of the final predictions being much less correlated with molecular weight. Moreover, our DL models and meta-models demonstrated their competitive power in comparison to other recently developed structure-based models in a large-scale affinity prediction benchmark and a virtual screening application benchmark.
@@
Overall, our framework corroborates existing work on ensemble learning, stacked generalization, super learner, and consensus-based improvements in predictions. Our machine-learning-based meta-models allow for the flexibility of adding or removing any component tools, physicochemical properties, or molecular descriptors as part of feature engineering. Models trained on multiple datasets (e.g. BindingDB and PDBbind) and on multiple data types (e.g. sequence and/or structure) may be unified to reduce individual biases and improve predictions. Furthermore, we posit that the DL models from DeepPurpose considered herein, being sequence-based, offer better scalability. Given that only ligand and protein sequences are required for the prediction, and not all-atom 3D structures, we expect that increasingly large databases could be used for training in future iterations by running high-throughput experimental binding affinity assays (without the need for concurrent X-ray or NMR experiments) as a filtering strategy in early drug discovery. Considering that the observed performances of these models for all our benchmarks are better or on par with more sophisticated structure-based DL models, there is an argument to be made in favor of more scalable approaches. Additionally, we note that the public and private experimental binding affinity databases are highly fragmented, in that they are produced separately, with potentially different types of assays, or at different scales. While efforts such as the PDBbind database manage to carefully curate a high-quality single dataset, not all data sources can be incorporated because of the type of data (e.g. experimental data without associated structures being available) or because of data access (e.g. behind a private paywall). Our meta-modeling may address some of these issues by pooling only prediction outputs from models trained on diverse public or private databases with or without available structures, protecting the databases themselves if necessary and maintaining performance quality. Subsequent in-depth analysis could also identify prediction synergy and complementarity of all available tools as we demonstrated in Fig. 6. Finally, there might be value in expanding families of meta-models with different input data, each family being optimized to target a particular protein or class of proteins and/or ligands as demonstrated in Tables 5 and 6.
@@
We acknowledge some limitations of our study. First, we primarily focused on a single dataset, PDBbind RefinedSet, for model training. In fact, in light of recent positive results from training models on the full PDBbind GeneralSet of ~19,000 complexes 42, 47, there is a scope for further scale-up of the model training process with larger data. It is worth mentioning, however, that when Jones et al. used the PDBbind general or refined set for model training, they observed the tradeoffs between data quality and quantity in performance of different models: SG-CNN performed better with the general set (in favor of data quantity); 3D-CNN performed better with the refined set (in favor of data quality). On the other hand, we started with DL models pre-trained on BindingDB of ~66,000 complexes and either employed them “out of the box” or fine-tuned them on the PDBbind RefinedSet (excluding the CoreSet). The two BindingDB-based meta-models, ED1 and ED2, perform better than any of the stand-alone docking or DL tools (Pearson correlation 0.62-0.64 vs. 0.47-0.59). We also tested the two PDBbind fine-tuned models, D1F and D2F, on the primary training set of BindingDB, to ascertain the degree to which the models retained their “out of the box” performance on the primary training set. We found reasonably good correlations for D1F and D2F (r = 0.636 to 0.711) (Supplementary Fig. S4) compared to the pre-trained D1 (r = 0.8) 57. However, the PDBbind-trained D3 showed poor prediction on BindingDB (r < 0.2). More importantly, while the CASF-2016 benchmark set has been well accepted, more benchmark sets of high quality would be needed to test and compare models in a more objective manner and also to challenge model generalization. Our second benchmark using the large-scale GeneralSet, although lower quality than CASF-2016, was such an effort to assess model generalization. Our virtual screening benchmark is also very limited, although our models outperformed the docking tools and tend to perform well for ligands with low molecular weight. In particular, the binary classification problem for screening of active or functional ligands would require a new model development for higher accuracy rather than repurposing exclusively affinity-based regression models as in this study. In this sense, the good performances of our fine-tuned DL models or meta-models do not imply model generalizability in an absolute sense. We also note that we neither fine-tuned the docking tools nor conducted hyperparameter optimization for the DL models, which might help performance as well.
@@
Our proof-of-concept study of meta-modeling can be readily extended to integrate with (i.e., to add more meta-features of) predictions from any other tools, such as HAC-Net, FAST, KDeep, and/or PerSpect ML, molecular properties or embeddings, or even predictions from the same base models trained on different orthogonal datasets, to address model bias and robustness, which is the key conceptual advantage of any ensemble or fusion model. In parallel to this type of horizontal extension (i.e., feature engineering), it is conceivable that meta-models may be expanded or stacked vertically, leading to deep meta-models. Our meta-modeling approach offers promising and flexible strategies for future model iterations to further improve ligand-protein binding affinity prediction.
@@
The BindingDB and PDBbind databases are publicly available. The code is partially available, due to our pending patent, on our GitHub page at https://github.com/Lee1701/Lee2023a.
@@ Acknowledgements
1Dept. of Genetics and 2Yale Center for Genome Analysis, Yale University, New Haven, CT 06510, USA
3Dept. of Molecular Biophysics & Biochemistry, Yale University, New Haven, CT 06520, USA
4Program in Computational Biology & Bioinformatics, 5Dept. of Computer Science, 6Dept. of Statistics & Data Science, and 7Dept. of Biomedical Informatics & Data Science, Yale University, New Haven, CT 06520, USA
@@
Preparation of the receptor structure: The receptor structure is read in as a PDB file and preprocessed in PyMol (https://pymol.org). We used the pymol class and associated functions within our pipeline script. The preprocessing consists of the removal of all solvent and heteroatoms in the molecules. The resulting PDB file is then converted into the requisite PDBQT format using OpenBabel (http://openbabel.org). We used the obabel flags ‘-e = continue after errors’, ‘-p = add hydrogens appropriate for the pH’, and ‘–partialcharge gasteiger’ to calculate Gasteiger partial charges. The pH was set at 7.0. We included the PDBQT-specific flags ‘r = output as a rigid molecular’ and ‘p = Preserve atom indices from input file’. The implementation of this is written as
@@
For the LIT-PCBA virtual screening exercise, we utilized sets of active and inactive ligands in the ratio of 1:2 randomly drawn from the combined training and validation cohorts of the LIT_PCBA database8 (see Methods). To run the docking tools on the datasets, we generated three-dimensional SDF structures for each of the ligands from their SMILES strings using the following command:
@@
[margin-bottom=0.0063in,margin-top=0mm,margin-right=0mm,margin-left=0mm,border-style=none,padding=0mm]![[Uncaptioned image]](extracted/5885633/a0000-img018.png)
@@
Figure S8. Performance of meta-models for PBDbind2020 GeneralSet as a secondary benchmark. Predictions by XGBoost with the Consensus filtering and no RMSD filter are shown.